OpenStudy (anonymous):

Differentiate the equation x^4+5xy^2=10 with respect to the variable t and express [dy/dt] in terms of [dx/dt]. (Use D for [dx/dt]). The variable "t" stuff confuses me greatly.

myininaya (myininaya):

\[[x(t)]^4+5x(t)[y(t)]^2=10\] ------------------------ \[[x(t)]^4=4[x(t)]^{4-1}x'(t)\] \[[5x(t)(y(t))^2]'=5[x(t)(y(t))^2]'=5[x'(t)(y(t))^2+x(t) \cdot 2(y(t))^{2-1}y'(t)]\] \[[10]'=0\]